### v2

@@ 3.3 Simulation Study
  • Model 1 accounts for a single fixed effect per school. Model 2 includes a single fixed effect per school with a dummy variable capturing isolated students. Model 3 is our structural model.
4  Empirical Illustration
In this section, we provide an empirical illustration of our econometric approach to estimating peer effects using a unique and now widely-used dataset from the National Longitudinal Study of Adolescent to Adult Health (Add Health). Specifically, our main objective in this section is to compare the estimate of peer effects obtained using our approach (equation 4) to that obtained using the classical linear-in-means peer effects specification (equation 5), in a context where isolated students are present.
4.1 Data
Table LABEL:tab:stat provides definitions of the variables included in the empirical analysis and presents summary statistics. The average GPA of students is 2.8, while the average GPA of their friends is 2.9. On average, students are approximately 15 years old and have attended their current school for 2.5 years. In terms of demographic composition, the sample includes 48.7% boys, 16.4% Hispanics, 64.7% whites, 16.8% Blacks, 7% Asians, and 9.5% from other racial backgrounds. Furthermore, 74.1% of the students live with both parents. Regarding maternal education, about 30.6% of students’ mothers have attained a high school (HS) education level, while 16.9% have education levels beyond HS, and 41.9% have education levels below HS. In terms of mothers’ occupations, 20.2% of the students’ mothers work in professional occupations (such as teachers, doctors, lawyers, and executives), 20% are homemakers or do not work, and 43.5% hold other jobs. [^14] Add Health provides a detailed list with more than 15 categories. We combine these occupations into four broader categories along with a missing indicator.
As an important feature of the proposed method is to account for whether students are isolated or not isolated, we also present summary statistics for the subsample of students who have no friends. The average GPA for this group is slightly lower than that of students who have friends. We also observe that students who have no peers are more likely to be males and are often from minority groups, such as Hispanics, Blacks, and Asians.

### v3

@@ Appendix E Supplementary Results on the Simulation Study and Application
  • Notes: Model 1 accounts for a single fixed effect per school. Model 2 includes a single fixed effect per school with a dummy variable capturing isolated students. Model 3 is our structural model.
Table 8: Detailed estimation results after excluding isolated students
Model 1 Model 2 Model 3 Model 1
Coef Sd Err Coef Sd Err Coef Sd Err Coef Sd Err
Peer Effects $0.561$ $0.030$ $0.788$ $0.042$ $0.878$ $0.044$ $0.581$ $0.034$
Has friends 2.449-2.449 $0.145$
Own effects
Female $0.182$ $0.006$ $0.165$ $0.007$ $0.165$ $0.007$ $0.186$ $0.007$
Age 0.008-0.008 $0.004$ 0.047-0.047 $0.004$ 0.045-0.045 $0.004$ 0.009-0.009 $0.004$
Hispanic 0.096-0.096 $0.011$ 0.094-0.094 $0.011$ 0.086-0.086 $0.011$ 0.092-0.092 $0.012$
Race
Black 0.113-0.113 $0.013$ 0.124-0.124 $0.013$ 0.102-0.102 $0.015$ 0.076-0.076 $0.016$
Asian $0.199$ $0.014$ $0.183$ $0.015$ $0.173$ $0.015$ $0.175$ $0.015$
Other 0.030-0.030 $0.011$ 0.032-0.032 $0.011$ 0.029-0.029 $0.012$ 0.023-0.023 $0.012$
Lives with both parents $0.098$ $0.008$ $0.088$ $0.008$ $0.083$ $0.008$ $0.098$ $0.008$
Years in school $0.032$ $0.003$ $0.025$ $0.003$ $0.023$ $0.003$ $0.027$ $0.003$
Member of a club $0.169$ $0.013$ $0.158$ $0.013$ $0.150$ $0.013$ $0.182$ $0.015$
Mother’s education
Less than HS 0.072-0.072 $0.009$ 0.066-0.066 $0.009$ 0.062-0.062 $0.009$ 0.072-0.072 $0.010$
More than HS $0.146$ $0.008$ $0.125$ $0.008$ $0.118$ $0.008$ $0.132$ $0.008$
Missing $0.017$ $0.013$ $0.012$ $0.013$ $0.013$ $0.013$ $0.007$ $0.014$
Mother’s job
Professional $0.040$ $0.009$ $0.037$ $0.010$ $0.034$ $0.010$ $0.029$ $0.010$
Other 0.035-0.035 $0.008$ 0.032-0.032 $0.008$ 0.031-0.031 $0.008$ 0.041-0.041 $0.008$
Missing 0.070-0.070 $0.012$ 0.064-0.064 $0.012$ 0.061-0.061 $0.012$ 0.078-0.078 $0.013$
Contextual effects
Female 0.122-0.122 $0.012$ 0.111-0.111 $0.013$ 0.127-0.127 $0.013$ 0.137-0.137 $0.012$
Age 0.082-0.082 $0.004$ $0.029$ $0.005$ $0.028$ $0.006$ 0.075-0.075 $0.005$
Hispanic $0.049$ $0.017$ $0.094$ $0.019$ $0.086$ $0.021$ $0.023$ $0.017$
Race
Black 0.010-0.010 $0.017$ $0.067$ $0.019$ $0.055$ $0.021$ 0.040-0.040 $0.019$
Asian 0.051-0.051 $0.022$ 0.111-0.111 $0.026$ 0.129-0.129 $0.028$ 0.052-0.052 $0.023$
Other 0.040-0.040 $0.020$ 0.008-0.008 $0.021$ 0.001-0.001 $0.022$ 0.028-0.028 $0.019$
Lives with both parents 0.047-0.047 $0.016$ 0.011-0.011 $0.017$ 0.021-0.021 $0.018$ 0.023-0.023 $0.016$
Years in school $0.032$ $0.004$ 0.010-0.010 $0.005$ 0.008-0.008 $0.006$ $0.039$ $0.005$
Member of a club 0.160-0.160 $0.028$ 0.065-0.065 $0.028$ 0.091-0.091 $0.029$ 0.180-0.180 $0.029$
Mother’s education
Less than HS 0.043-0.043 $0.016$ $0.020$ $0.018$ $0.026$ $0.019$ 0.022-0.022 $0.016$
More than HS $0.014$ $0.017$ 0.026-0.026 $0.020$ 0.036-0.036 $0.021$ $0.022$ $0.016$
Missing 0.068-0.068 $0.024$ 0.033-0.033 $0.026$ 0.031-0.031 $0.026$ 0.046-0.046 $0.023$
Mother’s job
Professional 0.062-0.062 $0.018$ 0.026-0.026 $0.019$ 0.036-0.036 $0.020$ 0.037-0.037 $0.017$
Other 0.103-0.103 $0.014$ 0.021-0.021 $0.016$ 0.021-0.021 $0.016$ 0.070-0.070 $0.014$
Missing 0.110-0.110 $0.021$ $0.004$ $0.024$ $0.010$ $0.024$ 0.089-0.089 $0.020$
$\sigma^{2}_{\eta}$ $0.296$ $0.292$
$\sigma^{2}_{\epsilon}$ $0.493$ $0.099$ $0.047$ $0.480$
$\rho$ $0.168$ $0.485$
Weak instrument F 160 100 105 112
Sargan test prob. 0.000 0.095 0.493 0.000
Number of schools 139 139 139 139
Number of students 61,183 61,183 61,183 53,529
  • Notes: Models 1, 2, and 3 are estimated using the subsample excluding fully isolated students (students who nominate no friends and who have not been nominated by others), whereas Model 2 is estimated using the sample excluding any isolated friends (students who nominate no friends). Models 1 and 1 control for unobserved school heterogeneity but do not include dummy variables capturing isolated students (i.e., $\alpha_{s}=0$ and $c_{s}$ varies across schools). Model 2 includes a single fixed effect per school with a dummy variable capturing isolated students (i.e., $\alpha_{s}$ may not be zero but is constant across schools, whereas $c_{s}$ varies). Model 3 is our structural model.
Table 9: Estimation results using a Tobit model
Model 1 Model 2 Model 3
Coef Sd Err Coef Sd Err Coef Sd Err
Peer Effects $0.224$ $0.019$ $0.378$ $0.018$ $0.480$ $0.018$
Has Friends 1.126-1.126 $0.092$
Own effects
Female $0.188$ $0.007$ $0.184$ $0.007$ $0.184$ $0.007$
Age 0.038-0.038 $0.004$ 0.048-0.048 $0.003$ 0.048-0.048 $0.004$
Hispanic 0.122-0.122 $0.011$ 0.122-0.122 $0.011$ 0.114-0.114 $0.011$
Race
Black 0.162-0.162 $0.013$ 0.167-0.167 $0.013$ 0.151-0.151 $0.014$
Asian $0.269$ $0.015$ $0.265$ $0.015$ $0.256$ $0.015$
Other 0.028-0.028 $0.012$ 0.028-0.028 $0.012$ 0.026-0.026 $0.012$
Live with both parents $0.126$ $0.008$ $0.125$ $0.008$ $0.122$ $0.008$
Year in school $0.037$ $0.003$ $0.036$ $0.003$ $0.034$ $0.003$
Member of a club $0.174$ $0.013$ $0.172$ $0.013$ $0.166$ $0.013$
Mother education
Less than HS 0.085-0.085 $0.010$ 0.083-0.083 $0.010$ 0.080-0.080 $0.010$
More than HS $0.188$ $0.008$ $0.182$ $0.008$ $0.178$ $0.008$
Missing $0.044$ $0.014$ $0.044$ $0.014$ $0.046$ $0.014$
Mother job
Professional $0.048$ $0.010$ $0.047$ $0.010$ $0.044$ $0.010$
Other 0.053-0.053 $0.009$ 0.052-0.052 $0.008$ 0.051-0.051 $0.008$
Missing 0.094-0.094 $0.012$ 0.093-0.093 $0.012$ 0.092-0.092 $0.012$
Contextual effects
Female 0.051-0.051 $0.014$ 0.065-0.065 $0.013$ 0.091-0.091 $0.013$
Age 0.044-0.044 $0.003$ 0.001-0.001 $0.005$ $0.004$ $0.006$
Hispanic $0.001$ $0.020$ $0.033$ $0.019$ $0.023$ $0.020$
Race
Black 0.048-0.048 $0.018$ $0.008$ $0.018$ $0.014$ $0.020$
Asian $0.043$ $0.025$ 0.002-0.002 $0.024$ 0.019-0.019 $0.025$
Other 0.048-0.048 $0.024$ 0.029-0.029 $0.023$ 0.016-0.016 $0.023$
Live with both parents $0.070$ $0.017$ $0.077$ $0.016$ $0.065$ $0.016$
Year in school $0.027$ $0.005$ $0.008$ $0.005$ $0.009$ $0.006$
Member of a club $0.043$ $0.028$ $0.085$ $0.027$ $0.065$ $0.026$
Mother education
Less than HS 0.065-0.065 $0.020$ 0.029-0.029 $0.020$ 0.020-0.020 $0.019$
More than HS $0.144$ $0.017$ $0.118$ $0.016$ $0.114$ $0.016$
Missing 0.018-0.018 $0.029$ 0.007-0.007 $0.028$ 0.007-0.007 $0.027$
Mother job
Professional $0.008$ $0.022$ $0.024$ $0.021$ $0.015$ $0.020$
Other 0.099-0.099 $0.017$ 0.058-0.058 $0.017$ 0.057-0.057 $0.017$
Missing 0.114-0.114 $0.025$ 0.058-0.058 $0.025$ 0.051-0.051 $0.025$
  • Notes: We reestimate Models 1, 2, and 3 in Table 3 by accounting for the bounded nature of the outcome.
[…]